the length of a rectangular playing field is 5 ft less than twice its width. if the perimeter of the playing field is 230 ft. find the length and width of the field.
Answer by scott8148(6628)

L=2W-5 ... 2L+2W=230 ... L+W=115 ... (2W-5)+W=115 ... 3W=120 ... W=40 ... L=75